OpenStudy (gabrielah96):

(1-3cosx-4cos^2x/sin^2x=1-4cosx/1-cosx

OpenStudy (loser66):

do what? prove? solve?

OpenStudy (gabrielah96):

prove

OpenStudy (loser66):

Hint first, ok? denominator : sin^2 = 1- cos^2 = (1-cos) \(\color{red}{(1+cos)}\) pay attention at the red part, we are gonna cancel it out with numerator

OpenStudy (loser66):

numerator: \(1 - 3cos-4cos^2 = cos^2 \color {blue}{+2cos}\) +1\(\color{blue}{-2cos}-3cos-5cos^2\) know why I do it? for the blue part, I add, then subtract, so nothing happen, right? for the black part, I just do -4cos^2 = cos^2 - 5cos^2, and put cos^2 at front and -5cos^2 at last. got me?

OpenStudy (loser66):

so far, the first 3 terms I put forms the form of (cos +1)^2 the last 3 terms becomes -5cos -5cos^2. I factor those terms by -5cos(cos +1)

OpenStudy (loser66):

now combine!!! \((cos +1)^2-5cos(cos +1)\) = \(\color{red}{(cos +1)}(cos +1 -5cos)\) =\(\color{red}{(cos+1)}(1-4cos)\) that's the numerator

OpenStudy (loser66):

combine the nume and the denominator , cancel the red part, you get the right hand side the proof done

OpenStudy (anonymous):

The left side can be factored as follows: Numerator: -4cos^2 x - 3cos x + 1 = (cos x + 1)(1 - 4cos x) Denominator: sin^2 x = 1 - cos^2 x = (1 - cos x)(1 + cox ) After simplification get the answer.